OpenStudy (anonymous):

The base of a triangle with height h and area A is given by the formula b=2A/h What is the base of a triangle with area 17 cm2 and height 5 cm?

OpenStudy (anonymous):

whats some of the answers

OpenStudy (anonymous):

i got that one what about this one!

OpenStudy (anonymous):

6.8

OpenStudy (anonymous):

b= 2A/h b= 2(17)/5 b=6.8

OpenStudy (anonymous):

The formula d = rt gives the distance traveled in time t at rate r. A delivery truck drives for 1.5 hours and travels 76 miles. What was the average speed of the truck?